OpenStudy (anitaasks):

Pendulum! At t=0, a 1m long pendulum is released after being displaced to the left side of the equilibrium position. At what time does the mass reach its lowest position?

OpenStudy (anonymous):

You can find the period of the pendulum first, then the time you are looking for is 1/4of the period \[ T=2\pi\sqrt{l/g}\]